Calculate the number of liters of solution which can be made from each of the following

3. A 2.0 M solution using 80.0 grams of NaOH

4. A 6.0 M solution using 126 grams of HNO 3
1
Expert's answer
2020-05-24T03:09:43-0400

Solution:

Molarity = Moles of solute / litres of solution

Therefore,

Litres of solution = Moles of solute / Molarity


1)

1 mol of NaOH has a mass of 40.0 g, so

Moles of NaOH = 80.0 g NaOH × (1 mol NaOH / 40.0 g NaOH) = 2.00 mol NaOH

Litres of solution = (2.00 mol) / (2.0 M) = 1.0 L

The volume of the solution is 1.0 L


Answer (1): 1.0 L of solution.


2)

1 mol HNO3 has a mass of 63.0 g, so

Moles of HNO3 = 126 g HNO3 × (1 mol HNO3 / 63.0 g HNO3) = 2.00 mol HNO3

Litres of solution = (2.00 mol) / (6.0 M) = 0.33 L

The volume of the solution is 0.33 L


Answer (2): 0.33 L of solution.